Simplify the following as much as possible leaved the answer in exact form:
sin 225 degrees ctn 300 degrees + sec 150 degrees cos 315 degrees

Answer by venugopalramana(3286)

sin 225 degrees ctn 300 degrees + sec 150 degrees cos 315 degrees
={SIN(180+45)}*{COT(360-60)}+{SEC(180-30)}*{COS(360-45)}
=-SIN(45)*{-COT(60)}+{-SEC(30)}*COS(45)
=(1/SQRT.2)((1/SQRT.3)-(2/SQRT.3)(1/SQRT.2)
=(1-2)/SQRT(3*2)=-1/SQRT.6=-SQRT.(6)/6